OpenStudy (anonymous):

I know how to solve simultaneous equations by graphing, but i do not know how to solve quadratic equations by grahing, can someone explain to me how i go about solving quadratic equations by graphing?

OpenStudy (tkhunny):

Add the equation y = 0 and solve simultaneously.

OpenStudy (anonymous):

actually,i was given a homework to draw the graph of f(x)=x2 +2x -3, then i was asked to use the graph to solve two quadratic equations. Does y=o have anything to do with helping me to get the solutions?

OpenStudy (anonymous):

by the way, how do iadd that equation y=o?

jimthompson5910 (jim_thompson5910):

in your y= menu where you type in the graphs, you add on 0 onto y2 so you will have this y1 = x^2 +2x - 3 y2 = 0

OpenStudy (anonymous):

i do not understand what to do. The two quadratic equations that i am to use the graph to solve are x2+2x-3=1.5 and x2-2x-3=-3

jimthompson5910 (jim_thompson5910):

you want to see where the two intersect right?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

then get everything to one side for both equations x^2+2x-3=1.5 x^2+2x-3-1.5 = 0 x^2+2x-4.5 = 0 -------------- x^2-2x-3=-3 x^2-2x-3+3 = 0 x^2-2x = 0 -------------- now type x^2+2x-4.5 for y1 and type x^2 - 2x for y2 so y1 = x^2+2x-4.5 y2 = x^2 - 2x

OpenStudy (anonymous):

Okay, i understand how you got those answers, but how am i to use the y1 and y2 to get the solutions?

jimthompson5910 (jim_thompson5910):

you're using a TI calculator right?

OpenStudy (anonymous):

no, i am not using any calculator

OpenStudy (tkhunny):

Maybe it would help if you provided an example of how you do this: "I know how to solve simultaneous equations by graphing," What equations can you solve by graphing?

jimthompson5910 (jim_thompson5910):

oh you're graphing this by hand?

OpenStudy (anonymous):

yes i am graphing it by hand

jimthompson5910 (jim_thompson5910):

if you can, I would use graphing tools to help you are you supposed to do it by hand and not use a calculator? or is it a case where you just don't have a calculator?

OpenStudy (anonymous):

i am to do it by hand

jimthompson5910 (jim_thompson5910):

if you have to do it by hand with no help from the computer, calculator, or other tools, then why not just use the quadratic formula?

OpenStudy (anonymous):

i drew my graph already but i do not know how to use the graph that i drew to solve the aforementioned pairs of quadratic equations

jimthompson5910 (jim_thompson5910):

oh you draw the two graphs on the same xy axis and see where they intersect

OpenStudy (tkhunny):

Did you put the equations on the same set of coordinate axes?

OpenStudy (anonymous):

Tkhunny i only drew one graph but i was asked to use the graph that i drew to solve the two pairs of quadratic equations.

OpenStudy (anonymous):

Jim thompson, there was not any need to draw two graphs

OpenStudy (anonymous):

the graph that i drew looks something like this